# Tiling Spaces and the Expanding Universe: Bridging Quantum Mechanics and Cosmology

### v1

@@ 3.1.1 Tiling Spaces: From Discrete to Continuous
This framework allows us to analyze the statistical properties of quasicrystalline structures and their implications for cosmology. The information entropy $I_{\varLambda}$ plays a crucial role in understanding the distribution and frequency of clusters within the tiling space, and the derived energy density connects this microscopic structure to macroscopic physical properties at large scales. With matter distributions, there are associated dispersion relations, which can lead to dynamics as we pointed out earlier.
Another aspect to consider is the self-similarity of quasicrystals. Quasicrystal point sets and associated tilings, while not retaining the discrete translational symmetry of the lattice in the projected space, exhibit new discrete symmetries such as scaling symmetry (self-similarity). This implies that the frequencies in a growing quasicrystal are dynamic and may oscillate.
Let us consider some suggestive interpretations of our working hypothesis of the universe as a growing quasicrystal.
3.1.2 Growth Dynamics in a Quasicrystalline Universe and Dark Energy
This model incorporates two distinct types of growth, each playing a crucial role in the universe’s evolution:
External Growth by Shell Projection: The first type of growth occurs through the projection of new points from a higher-dimensional lattice onto our observable universe. This process can be visualized as the universe growing outward, shell by shell, from a central point. Each new shell introduces additional points, expanding the universe’s structure.
Internal Growth through Subdivision and Rescaling: The second type of growth is an internal process triggered by the introduction of matter, which naturally imposes a scale on the system. This scale can be aligned with the Planck scale, introducing minimum length and time scales. As new points are projected into the universe, they subdivide existing structures. To maintain consistency with the matter-imposed scale, the universe must undergo periodic rescaling.
To illustrate these growth processes, let’s consider a simplified one-dimensional universe, the usual Fibonacci chain projected point by point:
Initial state: The universe begins with a single point and two tile lengths: short and long.
Second point: A new point arises. The original short tile becomes long, and a new short tile is introduced with the old long becoming a new long and a short.
Fourth point: With the addition of the fourth point, we now have three types of tiles. This illustrates the dynamic nature of the frequency distribution due to the asynchronicity between shell projections and the self-similarity requirements.
Fifth point: At this stage, we return to having two tile types, but at a different scale from the initial state.
If matter is defined based on the initial scale, the universe must rescale this new configuration to maintain consistency. The new long tile is rescaled to match the original long tile, and the same is done for the short tile. This rescaling process is crucial for maintaining the consistency of matter within the growing quasicrystal structure. It allows for the preservation of local minimum distances between points if the rescaling is precisely timed after a certain number of points is projected.
Matter and observers are ultimately composed of points that are continuously replaced by new points projected from a higher dimension. One analogy is the process in which a person’s atoms are replaced many times throughout their life. This mechanism for dark energy can be modeled by assuming an energy constraint, leading to $\Lambda$ in Equations 18 and 19.
The two distinct expansion processes may unfold differently in the early radiation-dominated epoch and the later matter-dominated epoch: Growth due to the construction of the quasicrystal shell by shell from projections from a higher dimension. Rescaling of internal subdivisions from points entering the quasicrystal from the projections, necessitated by matter fixing a scale. This dual growth mechanism qualitatively addresses the Hubble tension.
3.1.3 Phonons and Dark Matter
In our quasicrystalline universe model, we can extend the analogy with solid-state physics to propose a novel explanation for dark matter. This approach leverages the concept of phonons—quantized modes of vibrations in a quasicrystal structure—and applies it to the spacetime quasicrystal of our model.
In this framework, spacetime itself is viewed as a quasicrystal, and phonons represent vibrational modes within this spacetime quasicrystal. These phonons, particularly those with long wavelengths that persist to later times, could potentially contribute to the dark matter component of the universe. This concept draws a parallel with how phonons interact with electrons in solid-state physics. In solid-state systems, phonons can scatter electrons, influencing phenomena such as electrical resistance and superconductivity. The interaction Hamiltonian in solid-state physics typically takes the form:
$H_{\text{int}}=g\sum_{q,k}(a_{q}+a_{q}^{\dagger})c_{k+q}^{\dagger}c_{k},$ (39)
where $a_{q}$ and $a_{q}^{\dagger}$ are phonon annihilation and creation operators, respectively, and $c_{k}$ and $c_{k+q}^{\dagger}$ are electron operators.
Extending this concept to our cosmological model, we can propose an interaction Lagrangian that couples spacetime phonons with ordinary matter (e.g., electrons):
$\mathcal{L}_{\text{int}}=g_{ph}\sum_{q,k}(\phi_{q}+\phi_{q}^{\dagger})\psi_{k+q}^{\dagger}\psi_{k},$ (40)
where $\phi_{q}$ and $\phi_{q}^{\dagger}$ are the phonon fields, and $\psi_{k}$ and $\psi_{k+q}^{\dagger}$ are ordinary matter fields like electrons.
Dark Matter Candidate: These spacetime phonons, if weakly interacting, could serve as candidates for dark matter. They would influence the dynamics of ordinary matter while remaining largely undetectable through electromagnetic interactions. This is analogous to models where dark matter interacts with ordinary matter through new forces or mediators, modifying the dynamics of both dark matter and ordinary matter. For instance, similar to dark photon interactions in some models, phonons in the spacetime quasicrystal could potentially scatter off nuclei, producing recoil signals that might be detectable in experiments designed to look for Weakly Interacting Massive Particles (WIMPs). Furthermore, the decay or annihilation of these phonons might produce standard model particles that could be detected by telescopes or neutrino detectors. It’s important to note that the quantitative details of this phonon-based dark matter model are still a work in progress. However, this approach offers a novel perspective on dark matter, grounded in the quasicrystalline structure of spacetime proposed by our model. This phonon-based dark matter hypothesis not only provides a potential explanation for the nature of dark matter but also further illustrates the deep connections between solid-state physics and cosmology suggested by our quasicrystalline universe model.
3.2 Implications of the Expanding Universe as a Growing Quasicrystal

### v2

@@ 2.2 Tiling Space
This framework allows us to analyze the statistical properties of quasicrystalline structures and their implications for cosmology. The information entropy $I_{\varLambda}$ plays a crucial role in understanding the distribution and frequency of clusters within the tiling space, and the derived energy density connects this microscopic structure to macroscopic physical properties at large scales.
2.3 Phonons , Phasons, Effective Field Theory
Thus, the early growing quasicrystal universe can be described microscopically using quantum mechanics or the distribution and frequency of clusters within the tiling spaces as outlined above. However, when the universe becomes large enough to support phonons and phasons, the regime of large number of tiles mentioned above, a new phase emerge. When phonons and phasons are present, knowledge from solid state physics tell us that in the regime of short wavelengths, the dispersion relation becomes nonlinear, and the microscopic details of the lattice become important. Phonons become sensitive to the discrete nature of the lattice, and the nonlinear dispersion can lead to phonon dispersion, where phonon packets with different wavelengths propagate at different speeds, causing distortion of the waveform. However, the dispersion relation for phonons in lattices is similar to the relativistic dispersion relation in the regime of large wavelengths, but with the speed of light replaced by the speed of sound. In this regime, the phonon dispersion relation is nearly linear, and the crystal can be treated as a continuous elastic medium, with phonons behaving like classical sound waves. The transition between these two regimes occurs when the phonon wavelength becomes comparable to the interatomic spacing, with the specific wavevector or wavelength at which this transition happens depending on the details of the crystal structure and the interatomic forces.
In quasicrystals, the situation is more complex because they do not have a simple primitive cell like periodic crystals but rather a complex, self-similar structure. As a result, the phonon dispersion relations in quasicrystals can exhibit intricate, fractal-like patterns. In the context of cosmology, self-similar structures are reminiscent of fractal patterns, and there has been significant interest in exploring fractal approaches to general relativity and cosmology [19, 20, 21, 22]. These studies investigate the possibility that the distribution of matter in the universe exhibits fractal properties over a range of scales. By connecting our quasicrystal model to fractal cosmology, we align with previous efforts to describe the universe’s large-scale structure using self-similar patterns.
Despite the differences between phonons in crystals and quasicrystals, the linear dispersion relation for phonons in quasicrystals shares similarities with the phonon branches in periodic crystals, particularly at low frequencies and long wavelengths. Phasons in quasicrystals also share similarities with phenomena found in relativistic physics. However, a key difference is the presence of a dispersive imaginary term in the phason dispersion relation, which becomes prominent at very large wavelengths [23, 24]. This imaginary term is related to the diffusive behavior of phasons and is absent in the phonon dispersion relation.
We can extend our discussion by proposing an effective field theory (EFT) for quasicrystals that provides a concrete mechanism for setting the inflaton field, offering an alternative to standard reheating mechanisms. Following the approach in [23], we consider the phason field inherent in quasicrystals as a candidate for the inflaton.
In the regime of large wavelengths, the quasicrystal can be treated as a continuous elastic medium. This approximation allows us to model the quasicrystal using scalar fields within an EFT framework, aiming to reproduce the results of conventional inflationary models. As quasicrystals exist at finite temperatures, their equilibrium state is described by a thermal density matrix involving mixed states. Conventional quantum field theory techniques may not be fully appropriate in this context, necessitating the use of EFT methods tailored for finite-temperature systems.
2.3.1 Effective Lagrangian for Quasicrystals
We consider the effective action for quasicrystals in the classical limit, incorporating both phonon and phason degrees of freedom. The quasicrystal fields, denoted by $\psi^{A}$, represent internal rearrangements of the quasicrystal structure without displacing atoms. These fields are crucial as they can be interpreted as scalar fields in the EFT, making them ideal candidates for the inflaton field in cosmology. The effective Lagrangian for quasicrystals, as presented in [23], is given by:
$\mathcal{L}{\text{EFT}}=T^{\mu\nu}\partial\mu X_{a\nu}+J^{A\mu}\partial_{\mu}\psi_{a}^{A}+\Gamma^{A}\psi_{a}^{A}+\frac{i}{2}M^{AB}\psi_{a}^{A}\psi_{a}^{B}.$ (19)
Here:
  • $T^{\mu\nu}$ is the stress-energy tensor, capturing the energy and momentum densities of the system.
  • $J^{A\mu}$ are the currents associated with the quasicrystal fields, including both phonons and phasons.
  • $\Gamma^{A}$ and $M^{AB}$ are dissipative coefficients introducing non-equilibrium dynamics and dissipation into the system.
  • $X_{a\nu}$ and $\psi_{a}^{A}$ are the difference fields in the Schwinger-Keldysh formalism, defined as $X_{a\nu}=X_{1\nu}-X_{2\nu}$ and $\psi_{a}^{A}=\psi_{1}^{A}-\psi_{2}^{A}$, where the subscripts $1$ and $2$ denote the forward and backward branches of the contour, respectively.
The field $\psi^{4}$ will be interpreted as the phason field. It transforms under a shift symmetry $\psi^{4}\rightarrow\psi^{4}+\lambda^{4}$, where $\lambda^{4}$ is a constant vector in the higher-dimensional space characterizing the quasicrystal. This shift symmetry implies that the phonon field behave as scalar fields under translations in the quasicrystal’s higher-dimensional embedding space. Consequently, the phason field $\psi^{4}$ can be treated as scalar fields in the EFT framework, analogous to the inflaton field in cosmology.
2.3.2 Phason Field as the Inflaton
In our framework, the phason field $\psi^{4}$ acts as the inflaton field driving cosmic inflation. The choice of the component $\psi^{4}$ is motivated by its distinct role in the derivative expansion of the EFT, where its derivatives are counted differently compared to the spatial components $\psi^{i}$ ($i=1,2,3$).
The dynamics of $\psi^{4}$ are governed by the equations of motion derived from the effective Lagrangian (19). Focusing on the phason component, the equation of motion for the classical (or retarded) phason field $\psi_{r}^{4}=\frac{1}{2}(\psi_{1}^{4}+\psi_{2}^{4})$ is:
$H^{44}\partial_{\mu}\partial^{\mu}\psi_{r}^{4}+\frac{\partial F^{4}}{\partial Z^{4}}\partial_{0}^{2}\psi_{r}^{4}-M^{44}\partial_{0}\psi_{r}^{4}=0,$ (20)
where:
  • $H^{44}$ and $F^{4}$ are functions determined by the effective potential and kinetic terms in the Lagrangian.
  • $Z^{4}=\beta^{\mu}\partial_{\mu}\psi_{r}^{4}$, with $\beta^{\mu}$ being the inverse temperature four-vector.
  • $M^{44}$ is a dissipative coefficient associated with the phason field.
The equation of motion (20) describes the dynamics of a scalar field with dissipative terms, analogous to the inflaton field in cosmology. The presence of the dissipative coefficient $M^{44}$ introduces friction-like effects, which are significant in the context of warm inflation scenarios where the inflaton dissipates energy into radiation during inflation.
Transforming to Fourier space, we obtain the dispersion relation for the phason field:
$\omega^{2}+i\gamma\omega=v^{2}k^{2},$ (21)
where:
$\gamma=\frac{M^{44}}{H^{44}-\frac{\partial F^{4}}{\partial Z^{4}}},\quad v^{2}=\frac{H^{44}}{H^{44}-\frac{\partial F^{4}}{\partial Z^{4}}}.$ (22)
The dispersion relation (21) indicates that the phason field exhibits both propagating (real frequency) and dissipative (imaginary frequency) behavior. Using the relation $\omega=\dfrac{E}{\hbar}$ for the phason field, we rewrite the dispersion relation as:
$v^{2}k^{2}=\dfrac{E^{2}}{\hbar^{2}}+i\gamma\dfrac{E}{\hbar}.$ (23)
Here, $\gamma$ acts as the phason friction coefficient, representing the dissipative effects in the phason dynamics. This form explicitly shows how the energy $E$ of the phason relates to its momentum $k$, incorporating both the dispersive term $\dfrac{E^{2}}{\hbar^{2}}$ and the dissipative term $i\gamma\dfrac{E}{\hbar}$.
The phason field’s behavior as a scalar field with dissipative dynamics makes it an attractive candidate for the inflaton in cosmology. The shift symmetry of $\psi^{4}$ ensures that it can have a flat potential, a key requirement for slow-roll inflation. Additionally, the dissipative term allows the phason field to decay and transfer energy to matter fields during inflation, leading to particle production and reheating within the inflationary period.
2.3.3 Built-in Reheating Mechanism
The imaginary term $\frac{i}{2}M^{AB}\psi_{a}^{A}\psi_{a}^{B}$ in the Lagrangian introduces dissipation, which leads to particle production during the inflationary period. As the phason field evolves, it decays into matter and radiation fields through typical couplings present in the quasicrystal EFT. This process effectively reheats the universe without the need for a separate reheating phase, providing an alternative to standard reheating scenarios.
The decay rate of the phason field can be inferred from the dissipative coefficients $M^{AB}$ in the Lagrangian. The reheating temperature $T_{\text{reh}}$ can be estimated by considering the energy density transferred to radiation during the phason’s decay. This mechanism seamlessly transitions the universe from the inflationary phase to the radiation-dominated era, aligning with observational constraints on the reheating temperature.
2.3.4 Implications for Cosmic Inflation
By utilizing the phason field inherent in quasicrystals, we propose a novel approach to cosmic inflation that naturally incorporates dissipation and particle production. The unique properties of the phason field offer several advantages:
  • Scalar Field Dynamics: The phason field $\psi^{4}$ behaves as a scalar field in the EFT, with its dynamics governed by derivative interactions and shift symmetry. This aligns with the requirements for the inflaton field in cosmology.
  • Slow-Roll Inflation: The shift symmetry of the phason field ensures a flat effective potential, satisfying the slow-roll conditions necessary for a prolonged period of accelerated expansion.
  • Intrinsic Dissipation and Reheating: The dissipative terms in the effective Lagrangian lead to friction-like effects, causing the phason field to lose energy to particle production during inflation. This provides a built-in reheating mechanism that naturally transitions the universe to the hot, radiation-dominated phase.
  • Compatibility with Observations: The model can be tuned to produce a spectrum of primordial perturbations consistent with cosmic microwave background observations. The dissipative dynamics may also lead to unique signatures in the non-Gaussianities of the perturbations, offering potential observational tests of the model, a work in progress.
To review our working hypothesis that the universe is a growing quasicrystal, we propose an early quantum mechanical phase, followed by the natural emergence of phonons and phasons. Once the quasicrystal grows sufficiently large, such that the dynamics occur on distance scales much larger than the spacing between the underlying points, we expect the model to exhibit continuous symmetry in the tiling space. At this stage, effective field theory can be applied, with phonons and phasons modeled as scalar fields. The interplay between phonon and phason modes may contribute to the generation of mass scales through mechanisms akin to dynamical symmetry breaking, where collective excitations lead to effective masses for particles.
2.4 Phonons and Dark Matter
In our quasicrystalline universe model, we extend the analogy with solid-state physics to propose a novel explanation for dark matter. The phonon fields $\psi^{i}$ in Equation (19) correspond to collective excitations (vibrations) of the quasicrystal lattice in spacetime. Unlike phasons, these fields follow conventional dispersion relations and could survive as relics, interacting weakly with ordinary matter and potentially contributing to the dark matter component of the universe.
Much like axions [25, 26, 27], which are not thermally produced but arise as classical oscillations of the axion field, giving rise to cold dark matter, phonons in our model emerge from the quasicrystal field through an EFT framework. Both axions and phonons are characterized by long wavelengths, enabling them to contribute to the dark matter component of the universe while remaining difficult to detect through electromagnetic interactions.
Additionally, just as WIMPs are theorized to interact weakly with ordinary matter, phonons in the spacetime quasicrystal could potentially scatter off nuclei, producing recoil signals detectable in experiments designed to search for WIMPs. The weak interaction between spacetime phonons and ordinary matter presents a similar observational challenge to that of axions and WIMPs, offering an alternative mechanism for dark matter.
By drawing on these analogies, we propose that spacetime phonons could serve as dark matter candidates. They would influence the dynamics of ordinary matter while remaining largely undetectable, providing a novel perspective grounded in the quasicrystalline structure of spacetime proposed by our model.
2.5 Growth Dynamics in a Quasicrystalline Universe and Dark Energy
Our quasicrystalline universe model incorporates two key growth mechanisms:
External Growth by Shell Projection
The universe expands externally through the projection of new points from a higher-dimensional lattice onto our observable space. Specifically, we consider the projection of the 8-dimensional $E_{8}$ lattice onto 4-dimensional space to form the ESQC. This process adds concentric shells of points, effectively growing the quasicrystal—and thus the universe—shell by shell.
Internal Growth through Subdivision and Rescaling
As new points are projected, they may appear within existing structures, leading to internal subdivision. Due to the quasicrystal’s self-similarity, the tile structures are preserved at different scales. To maintain consistency with a minimal length scale imposed by matter (e.g., the Planck scale), the universe undergoes rescaling. Whenever new points introduce distances smaller than this minimal scale, a global rescaling adjusts all structures to preserve fundamental physical constants.
This mechanism can be more easily grasped in lower-dimensional quasicrystals, such as the Fibonacci chain (one-dimensional) and the Penrose tiling (two-dimensional), which can be built with just two tile types. Steinhardt’s one-tile quasicrystal [28] further demonstrates how tiles keep dividing but reappear at different scales due to self-similarity. In three dimensions, quasicrystals like the 3D Penrose tiling involve more complex arrangements but follow the same principle [29]. In four dimensions, with the ESQC, the number of tiles is not well known, but the building block is the 600-cell polytope, which keeps arising at different scales.
Whenever a new shell introduces a shorter minimum distance, a global rescaling is applied to the entire quasicrystal. This rescaling adjusts all distances to ensure consistency with the previously established minimum distance, preventing distances from shrinking below a critical threshold and ensuring that the quasicrystal maintains a fixed internal structure.
Figure 3: The image demonstrates the stepwise increase in distance between two points in the ESQC, with rescaling applied when necessary. We project the first 10 shells of the E8 lattice one by one into 4D space. The distance from the origin in the first shell is fixed as the minimum reference distance. As subsequent shells are projected, new points may introduce shorter distances, and whenever this occurs, a global rescaling is applied to maintain the initial minimum distance. The linear progression in the image reflects the cumulative effect of this rescaling mechanism as the quasicrystal grows.
Implications for Dark Energy
The rescaling process requires energy to adjust the quasicrystalline structure while maintaining consistency with physical laws consistent. This energy can be associated with an effective dark energy component, contributing to the cosmological constant $\Lambda$ in the Friedmann equations. The continuous introduction of new points and the accompanying rescaling provide a natural mechanism for dark energy arising from the quasicrystalline structure of spacetime.
Addressing the Hubble Tension
The dual growth mechanisms may impact cosmological observations differently over time:
  • Early Universe: External growth dominates, with rapid shell projections expanding the universe.
  • Later Universe: Internal growth and rescaling become significant as matter influences the minimal scale, potentially affecting the observed expansion rate.
This interplay offers a qualitative framework to address the Hubble tension, the discrepancy between early and late-time measurements of the Hubble constant, by suggesting that variations in the expansion rate are tied to the quasicrystalline growth dynamics.
3 From Schrödinger to Friedmann Equation
In the previous sections, we proposed a heuristic model where the cosmic expansion observed at large scales emerges intrinsically from the growth dynamics of a quasicrystalline structure underlying spacetime. We considered how this quasicrystalline universe grows both externally, through the projection of new points from higher dimensions, and internally, via subdivision and rescaling to maintain consistency with a minimal length scale. Now, we aim to address how this theoretical growth, even at small scales, might affect quantum matter.
$G_{\mu\nu}+\Lambda g_{\mu\nu}=\frac{8\pi G}{c^{4}}T_{\mu\nu},$ (24)
where $G_{\mu\nu}$ is the Einstein tensor representing the curvature of spacetime, $g_{\mu\nu}$ is the metric tensor describing the geometry of spacetime, $T_{\mu\nu}$ is the stress–energy tensor reflecting the distribution and flow of energy and momentum in spacetime, $\Lambda$ is the cosmological constant, $G$ is the Newtonian constant of gravitation, and $c$ is the speed of light in vacuum.
On the other hand, quantum mechanics introduces the principle of wave-particle duality, which manifests both wave-like and particle-like properties in matter and energy, as illustrated by the De Broglie relations and the Schrödinger equation-Equations (25, 26).
Notably, both General Relativity and Quantum Mechanics relate matter to an underlying structure—geometric in the case of spacetime curvature, and wave-like in the case of quantum wavefunctions. This observation motivates us to explore a simplified model within the cosmological context, where the wave aspect of quantum mechanics is proportionally linked to the geometry of spacetime.
The Friedmann equations, which are key to understanding the universe’s expansion, are derived from general relativity under the assumption of a homogeneous and isotropic universe. These equations describe how the scale factor a(t) evolves over time due to the energy content of the universe. Our aim is to investigate whether analogous dynamics could emerge from the effects of spacetime rescaling on local quantum states, focusing on the later evolution in flat space at small quantum scales. The early universe’s curvature’s regime will be addressed in a separate paper.
Conventionally, the expansion of the universe is significant at large scales, particularly in the regions between galaxies, while gravitational forces within galaxies and smaller structures prevent any noticeable effect. However, this does not rule out the possibility that spacetime expansion occurs even within matter-dominated regions, albeit in ways that are difficult to observe directly. To explore this, we examine how a quantum state behaves in an expanding universe.
Here we consider the flat space scenario. The time-independent Schrödinger Equation for the wave function (energy eigenfunction) $\Psi(x)$, with $x$ representing $(x,y,z)$, of a particle with mass $m$ under potential $V_{QM}(x)$. The equation is given by:
$\nabla^{2}\Psi=-\frac{2m(E-V_{QM})}{\hbar^{2}}\Psi\quad,$ (25)
For a particle under a potential $V_{QM}(x)$, the time-independent Schrödinger equation can yield different forms of solutions depending on the relationship between the energy $E$ of the particle and the potential $V_{QM}$. We consider the particle in a box situation with a finite potential inside and infinite outside. The box may be taken to be spherical with a given radius. Specifically, inside the box, for a constant potential, being zero, or when $E>V_{QM}$, the solutions are of the form of plane waves. Conversely, for $E<V_{QM}$, the solutions differ significantly. When these solutions are substituted back into the time-independent Schrödinger equation, we obtain:
$k^{2}=\frac{2m(E-V_{QM})}{\hbar^{2}},$ (26)
where $k$ is the wavenumber.
$\eta(t_{2})=\frac{a(t_{2})^{3}}{a(t_{1})^{3}}\eta(t_{1}),$ (27)
Introducing time is essential to account for the rescaling effects. The problem now should be addressed by the time-dependent Schrödinger equation, as the potential depends on the dynamic radius or volume. This problem has been studied using various approaches [30] and references therein. In this context, the energies become dynamic, and a notion of energy density in quantum mechanics arises naturally [31]. Crucially, the wavenumber, wavelength, and quantum amplitudes become functions of the dynamical size of the box, and we posit that they are subject to the Hubble-Lemaître’s law explicitly. By connecting the geometry with the wave aspect of the particle-wave duality mentioned earlier, we aim to derive Friedmann-like equations. While Hubble-Lemaître’s law is empirically validated at cosmological distances, we extend its conceptual framework to ’intermediate’ or even microscopic scales as a theoretical proposition. These scales are not conventionally associated with cosmic expansion, but within our growing quasicrystal model, the underlying spacetime structure allows for such an extension. This hypothesis serves as a foundation to explore the consequences of spacetime expansion on quantum states confined within matter-dominated regions.
Revisiting equation (26), we modify it by dividing both sides by the dynamical volume $\eta$, which confines the particle in space. Equation (26) encapsulates the wave-particle duality, highlighting that the matter wavelength, $\lambda=\frac{2\pi}{k}$, is a fundamental quantum mechanical attribute that determines the probability density of locating the particle within the configuration space. Notably, a particle’s wavelength, inversely related to its momentum (and thus energy), serves as a bridge to cosmological dynamics. By postulating the Hubble-Lemaître’s law at small scales, expressed as $v=H\lambda$, where $H=\frac{\dot{a}}{a}$ is the time-dependent Hubble parameter, we correlate the particle’s wavelength to the cosmic scale factor $a(t)$ while keeping the energy constant. In our analysis, $\lambda$ represents the instantaneous wavelength of the particle’s wavefunction within the expanding box. It is the ‘local’ wavelength affected by the expansion of the confinement volume, rather than an observed wavelength subject to cosmological redshift. Since we are modeling the particle within a dynamically expanding potential, $\lambda$ changes as a function of time due to the expansion, reflecting how the particle’s quantum state adapts to the changing spatial boundaries. We focus on this local effect rather than the observed redshift because we are interested in how the expansion influences the quantum properties of the particle itself, not how an external observer perceives changes in wavelength over cosmological distances. This association suggests an emergent effective kinetic energy tied to the velocity vv associated with the expansion. We acknowledge that, in standard cosmology, recession velocities resulting from the expansion of space are not associated with kinetic energy in the Newtonian sense, as they do not correspond to motion through space but rather to the metric expansion of space itself. As discussed by Sharma [32], kinetic energy calculations involving galaxies must consider both peculiar velocities (which contribute to kinetic energy) and recessional velocities due to cosmic expansion (which do not contribute to kinetic energy in the traditional sense). However, within our theoretical framework, we consider the impact of the expansion on a localized quantum system confined within an expanding potential. Here, we introduce the effective kinetic energy as a conceptual parameter tied to the expanding volume, representing the energy associated with the expansion-induced changes in the quantum system’s boundary conditions. This effective kinetic energy is not due to the particle moving through space in the classical sense but emerges from the changing spatial confinement resulting from the expansion of the universe.
Consequently, equation (26) is transformed into
where $K=\frac{1}{2}mv^{2}$, $\rho=E/\eta$ represents the energy density, and setting $V_{QM}=0$. Note that we are simplifying the concept of energy density here. The particle energy is time-dependent, and any particle would be subject to the same expansion that occurs everywhere. This brings a notion of a particle horizon and an energy density within each particle volume expansion. This equation bears a striking resemblance to the first Friedmann equation, suggesting that the dynamics of a quantum particle in a box with time-dependent boundaries may provide insights into the expansion of the universe by establishing a connection between the particle’s wave nature and the cosmic scale factor. By using the effective kinetic energy as a conceptual tool, we aim to explore the potential interplay between quantum mechanics and cosmology within our model, acknowledging that this is a theoretical extension and not empirically verified at small scales. To align with the usual first Friedmann equation’s structure, the term $\frac{K\eta}{\pi^{2}\hbar^{2}}$ must remain constant
$K\eta=\frac{8\pi^{3}\hbar^{2}G}{3c^{2}}$ (29)
$H^{2}=\left(\frac{\dot{a}}{a}\right)^{2}=\frac{8\pi G}{3c^{2}}\rho,$ (30)
$(\hbar kc)^{2}=E^{2}-(mc^{2})^{2}.$ (31)
$C=E-K-V=0,$ (32)
where $E$ is the total energy, $K$ is a effective kinetic energy, and $V$ is a effective potential energy, all associated with the dynamical expansion process. This constraint allows us to examine the system’s properties across different scales, connecting the microscopic energy $E$ to macroscopic scales potentially represented by $K$ and $V$. Here, $K$ and $V$ can be viewed as effective parameters encapsulating average or microscopic system properties as the system expands.
The energy-balancing constraint in Equation (32) provides a framework for understanding how the total energy $E$ can be viewed in the expansion side as partitioned into kinetic and potential components during the expansion process. By considering the conservation of energy and the interplay between these different forms of energy, we can gain insights into the dynamics of the system across various scales. In the context of the relativistic particle in a box with moving walls, the application of this constraint may lead to the emergence of terms analogous to the energy density $\rho$, pressure $p$, and possibly the cosmological constant $\Lambda$ when deriving the relativistic Friedmann-like equation. By incorporating the energy-balancing constraint into the analysis, we can explore how the quantum mechanical properties of the system, such as the wave nature of the particle and the dynamical evolution of the wavefunction, relate to the classical cosmological parameters that describe the expansion of the universe.
Isolating $K$ while keeping $E$ and $V$ independent, we obtain
$E^{2}=K(E+V+V^{2}/K).$ (33)
Since the wave number is inversely proportional to the wavelength, we can apply the Hubble-Lemaître’s law as done before. Then, Equation (31) can be rewritten as:
with $\beta=v/c$, $\rho=E/\eta$ is the energy density, and $p=V/\eta$ is the pressure. Equation (34) is analog of the Friedmann equation for relativistic matter, derived from the quantum mechanical description of a particle in a box with moving walls. The first term on the right-hand side resembles the classical Friedmann equation. The second term, proportional to $V^{2}$, may be interpreted as related to the cosmological constant $\Lambda$ as it does not explicitly depend on $a$.
To obtain the usual contribution to radiation, we set $mc^{2}=0$. Typically, radiation energy density enters the Friedmann equation as proportional to $1/a^{4}$, while non-relativistic matter is proportional to $1/a^{3}$. Here, we derive a notion of energy density for individual massive particles and photons subject to dynamical boundary conditions related to the expansion of space, and they share a similar form. The $1/a^{4}$ factor for radiation arises from considering the number density of photons, which scales like matter, times the photon energy, which scales with $1/a$ due to redshift as the photon wavelength expands. The redshift here is accounted for through the Hubble-Lemaître’s law applied to the wavelength.
Importantly, we note that different dispersion relations will lead to different dynamics for the scale parameter, and various relations may be considered. This is in line with the usual approach where different cosmological epochs have different evolutions due to varying energy densities.
Let us visit now the dispersion relation for phonons, Equation (23). In the case of lattices, the energy states are defined in a precise polytope of the lattice, the Voronoi cell. For quasicrystals, there are more fundamental cells or tiles, but with one fixed, we can consider the expansion of that cell and, as before, apply the Hubble-Lemaître’s law to the wavelength. We also need to take in account the different regimes of short and large wavelength. To deal with the imaginary part, this equation can be interpreted from the energy relation where in the regime of short wavelength the phason propagation dominates with $v^{2}k^{2}=\frac{E^{2}}{\hbar^{2}}$ leading to a Friedmann like equation similar to Equation (34) as we can use Equation (33) and multiply by the volume on both sides of Equation (23), yielding
$H^{2}=\frac{K\eta}{4\pi^{2}\hbar^{2}}(\rho+p)+\frac{1}{4\pi^{2}\hbar^{2}}V^{2}$ (35)
where we considered the velocity associated to phasons equal to the one of the expansion.
In the regime of large wavelength we have $v^{2}k^{2}=i\gamma\frac{E}{\hbar}$ which leads to diffusion and decoupling from expansion contribution
$H^{2}=i\gamma\frac{E}{4\pi^{2}\hbar}$ (36)
$\dot{\rho}+3\frac{\dot{a}}{a}(\rho+p)=0.$ (37)
Taking the derivative in Equations (28), (34), and (35), with $V$ constant and $\beta=1$, we get for Equations (28), (34) and (35):
$2H\dot{H}=\alpha\dot{\rho},$ (38)
with $\alpha=K\eta/\pi^{2}\hbar^{2}$. Now, we can derive the acceleration term in both situations. For Equation (28), substituting the continuity equation into the derivative of the Friedmann-like equation yields (with the original potential equal to zero)
$\frac{\ddot{a}}{a}=-\frac{\alpha}{2}(\rho+3p).$ (39)
For Equations (34), with $m=0$
$\frac{\ddot{a}}{a}=-\frac{\alpha}{2}(\rho+p)+\Lambda$ (40)
with $\Lambda=\frac{1}{\pi^{2}\hbar^{2}}V^{2}$.
For Equation (35), the acceleration equation becomes
$\frac{\ddot{a}}{a}=-\frac{\alpha}{4}(\rho+p)+\frac{1}{4}\Lambda$ (41)
These acceleration equations provide insights into the dynamics of the scale factor in the presence of matter, radiation, and phason contributions. They demonstrate how the quantum mechanical description of particles in a box with moving walls can give rise to cosmological equations that resemble the Friedmann equations of general relativity. The inclusion of the phason term highlights the potential role of quasicrystalline structures and their associated quasiparticles in the evolution of the universe.
Our analysis commenced with the time-independent Schrödinger equation for a particle confined within a potential well. We then introduced several crucial modifications to the system: a general energy-balancing constraint Equation (32), a time-dependent boundary condition representing moving walls, and a dynamically evolving volume. These additions were aimed at establishing a connection between the quantum and classical regimes across different scales. By applying the Hubble-Lemaître’s law to the wavelength of the particle and relating it to the cosmic scale factor, we derived a Friedmann-like equation Equation (28) from the quantum mechanical description of the system.
In the relativistic extension of our analysis, we employed the Klein-Gordon equation and the relativistic energy-momentum relation Equation (31). This led to the emergence of a new term in the resulting Friedmann-like equation Equation (34), which bears a resemblance to the cosmological constant term in the standard Friedmann equations of general relativity.
Furthermore, we explored the implications of different dispersion relations, particularly those associated with phonons and phasons in quasicrystalline structures. The presence of an imaginary term in the phason dispersion relation Equation (23) is interpreted as having different regimes contributions.
In summary, our analysis bridges the quasicrystal model of the universe with quantum mechanical systems, showing how the underlying expanding quasicrystalline structure affects the behavior of quantum matter. We demonstrated that, within this framework, there exists a duality between quantum states, seen either as wave-like matter or as an expression of the evolving geometry through the scale parameter. This duality draws a parallel with Einstein’s vision, where matter and geometry are deeply interconnected. By relating the matter wavelength to the expansion of space, we linked the Schrödinger equation to the Friedmann equation. This suggests that the classical world’s emergence may be intrinsically tied to the expanding substrate, where we perceive either matter or expansion, but not both simultaneously at local scales. The quantum mechanical description naturally adapts to this evolving background, providing new insights into the relationship between quantum states and cosmic expansion. We also note that the phonon contribution in the early short-wavelength regime introduces a term, $1/4\Lambda$, which may offer a potential resolution to the Hubble tension, another workin in progress.
4 Discussions, Exploratory Insights and Future Perspectives
4.1 Implications of the Universe as a Growing Quasicrystal